# Video LLMs for Temporal Reasoning in Long Videos

### v3

@@ Abstract
This paper introduces TemporalVLM, a video large language model (video LLM) capable of effective temporal reasoning and fine-grained understanding in long videos. At the core, our approach includes a visual encoder for mapping a long-term input video into features which are time-aware and contain both local and global cues. In particular, it first divides the input video into short-term clips, which are jointly encoded with their timestamps into time-sensitive local features. Next, the local features are passed through a bidirectional long short-term memory (BiLSTM) module for global feature aggregation. The extracted time-aware and multi-level features are important for accurate temporal reasoning and fine-grained understanding in long videos. Moreover, to facilitate the evaluation of TemporalVLM, we present a large-scale long video dataset of industry assembly processes, namely IndustryASM, which consists of videos recorded on factory floors with actions and timestamps annotated by industrial engineers for time and motion studies and temporal action segmentation evaluation. Finally, extensive experiments on datasets of long videos, including TimeIT and IndustryASM, show that TemporalVLM achieves superior performance than previous methods across temporal reasoning and fine-grained understanding tasks, namely dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation. To the best of our knowledge, our work is the first to incorporate LSTMs into video LLMs.
@@ 1 Introduction
In this paper, we propose TemporalVLM, a video LLM capable of effective temporal reasoning and fine-grained understanding in long videos. Fig. 1 shows the differences in model architectures of TemporalVLM and previous works. Our model includes two key architectural contributions for temporal reasoning in long videos. Firstly, we divide a long-term input video into multiple short-term clips and introduce a time-sensitive clip encoder for extracting time-aware local features from each clip. Secondly, we adopt a bidirectional long short-term memory (BiLSTM) module which takes all the local features as inputs and computes global features from multiple clips. As a consequence, our features not only are time-sensitive but also contain both local fine-grained and global semantic information, which are crucial for temporal reasoning in long videos. Moreover, to further evaluate our model, we present IndustryASM, a large-scale long video dataset of industry assembly processes for temporal action segmentation benchmarking and time and motion studies. Our IndustryASM dataset comprises of $4803$ videos with an average video duration of $119$ seconds. It covers in total $52$ diverse industry assembly tasks and includes timestamp and action labels. We convert the labels into chat samples with manually written instructions. To the best of our knowledge, it is the first large-scale video dataset recorded on factory floors and labeled by industrial engineers for temporal action segmentation evaluation and time and motion analyses. Lastly, extensive experiments on datasets of long videos, i.e., TimeIT [35] and IndustryASM, demonstrate that our TemporalVLM approach achieves superior results over previous methods across temporal reasoning and fine-grained understanding tasks, namely dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation.
@@
  • We develop TemporalVLM, a video LLM for temporal reasoning and fine-grained understanding in long videos. Given a long-term input video, we first sample short-term clips and encode each clip along with its timestamps into time-aware local features. The local features are then fed to a BiLSTM module to extract global features. Therefore, our features not only are time-sensitive but also capture effective local and global information.
@@ 3 TemporalVLM: A Video LLM for Temporal Reasoning in Long Videos
Figure 2: Given a long-term input video with timestamps, we divide it into short-term clips with timestamps. The time-aware clip encoder is then employed on each clip along with its timestamps for extracting time-aware local features, which capture local fine-grained cues within each clip. Next, the bidirectional long short-term memory module is adopted for computing global features, which capture global temporal relationships across multiple clips. Lastly, the large language model takes video and text tokens as inputs and generates responses.
@@ 3.1 Time-Aware Clip Encoder
We propose a time-aware clip encoder (i.e., blue block in Fig. 2) for extracting local fine-grained cues within a clip. Given a long-term input video, we first decompose it into $C$ short-term clips and uniformly sample a fixed number of frames $N^{c}_{f}$ from each clip. Here, we set $C=6$ clips and $N^{c}_{f}=96$ frames per clip. The sampled frames for each clip along with their corresponding timestamps are then passed to our time-aware clip encoder for simultaneously encoding the frame contents and their timestamps, yielding time-aware local features. We adopt the time-aware frame encoder and sliding video Q-Former of TimeChat [35] as our time-aware clip encoder. The time-aware frame encoder uses a pre-trained image encoder [39] to obtain frame features, which are jointly encoded with timestamps via an image Q-Former [5], yielding time-aware frame features. These frame features are then fed to a sliding video Q-Former [35] for temporal modeling, yielding time-aware clip features for each clip. Recent works, e.g., LongVLM [49], leverage clip sampler and clip encoder for long video understanding but their encoders are not time-aware and hence are not effective for temporal reasoning.
@@ 3.2 Bidirectional Long Short-Term Memory
We introduce a BiLSTM module (i.e., green block in Fig. 2) for computing global features across multiple clips. We first concatenate time-aware local features extracted from clips in the temporal order which the clips appear in the video. We then pass the sequence of local features to a BiLSTM module for aggregating global features. Our BiLSTM module follows a standard architecture, including two LSTM networks: one processes the sequence in the original order (forward LSTM) and another processes the sequence in the reverse order (backward LSTM) as:
@@
where $\mathbf{h}^{f}_{t}$ and $\mathbf{h}^{b}_{t}$ denote the hidden states at time step $t$ of the forward LSTM and backward LSTM respectively and $\mathbf{c}_{t}$ is the input at time step $t$. The final output $\mathbf{h}_{t}$ at time step $t$ is obtained by concatenating the outputs of the forward LSTM and backward LSTM as:
$\mathbf{h}_{t}=[\mathbf{h}^{f}_{t},\mathbf{h}^{b}_{t}].$ (3)
Please refer to the textbook [18] for a detailed description. Our BiLSTM module utilizes both past and future information and is capable of capturing long-range temporal relationships in both forward and backward directions. As empirically observed in Sec. 5.3, BiLSTM outperforms various alternatives, including average pooling, linear layer, traditional LSTM, and transformer [43]. Our final features output by BiLSTM not only are time-sensitive but also contain both local fine-grained and global semantic information, which are important for temporal reasoning in long videos. Unlike the recent LongVLM model [49], which applies pooling operation to obtain global features, our BiLSTM module has learnable parameters and achieves better results as shown in Sec. 5.
@@ 4 IndustryASM: A Large-Scale Long Video Dataset of Industry Assembly Processes
We follow prior works [32, 13, 14, 35] to test TemporalVLM on dense video captioning, temporal video grounding, and video highlight detection. To further evaluate our model, we introduce IndustryASM, a large-scale long video dataset of industry assembly processes for temporal action segmentation and time and motion studies. To our best knowledge, it is the first large-scale video dataset recorded on factory floors and labeled by industrial engineers. Below we provide some details and annotation process.
@@ 5 Experiments
Implementation Details. We follow TimeChat [35] to use LLaMA-2 7B [42] as the LLM, and ViT-G/14 from EVA-CLIP [39] as the image encoder. We take InstructBLIP’s checkpoints to initialize the image Q-Former, and Video-LLaMA’s [56] checkpoints to initialize the video Q-Former. The BiLSTM module includes a forward LSTM with one hidden layer and a backward LSTM with one hidden layer. The parameters of the BiLSTM module and projection layers are initialized randomly. The LLM is fine-tuned using LoRA [12] with rank 32. We use 6 clips and sample 96 frames from each clip. We fine-tune our model on a subset of TimeIT [35] and Valley [27] for 7 epochs on a single 8-A100 (80 GB) machine. The parameters of the LLM and image encoder remain frozen throughout the fine-tuning. Please see our supplementary material for more details.
@@ 5.1 Zero-Shot Results
We first evaluate all methods in the zero-shot setting and present the results in Tab. 1.
5.1.1 Dense Video Captioning Results
For dense video captioning, it can be seen from Tab. 1 that our TemporalVLM model achieves the best results across all three metrics, i.e., SODA_c, CIDEr, and F1 Score. For example, we outperform TimeChat [35] by +0.3 on CIDEr and +0.5 on F1 Score. This is likely because TimeChat [35] treats the entire video as a single clip and aggregates video tokens via query aggregation, overlooking fine-grained information. In contrast, our extracted features, which capture both local fine-grained cues from each clip and global temporal dependencies across multiple clips, are effective for caption generation and event localization. Moreover, our approach obtains superior results than LongVLM [49] across SODA_c (+0.4), CIDEr (+1.2), and F1 Score (+0.8), which confirms the importance of our time-aware clip encoder and BiLSTM for dense video captioning.
5.1.2 Video Highlight Detection Results
The video highlight detection results in Tab. 1 show that our TemporalVLM model obtains the best performance on both mAP and HIT@1. In particular, it surpasses TimeChat [35] by +1.9 in mAP and +7.4 in HIT@1, while outperforming LongVLM [49] by +5.3 at mAP and by +16.3 at HIT@1. These results suggest the effectiveness of our model for video highlight detection. By aggregating global features from time-aware local features, it identifies the most relevant timestamps and assigns them saliency scores based on their ranks of importance within the video against a given query. Though LongVLM [49] extracts multi-level features, the lack of a time-aware encoder hinders their results.
5.1.3 Temporal Video Grounding Results
It is evident from Tab. 1 that TemporalVLM achieves the best performance for temporal video grounding across R@1 (IoU=0.5) and R@1 (IoU=0.7). Specifically, it outperforms TimeChat [35] by +2.2 and +0.9 on R@1 (IoU=0.5) and R@1 (IoU=0.7) respectively, and LongVLM [49] by +16.2 and +7.1 on R@1 (IoU=0.5) and R@1 (IoU=0.7) respectively. The results validate the benefits of our time-aware and multi-level features for temporal video grounding. Our time-aware clip encoder extracts fine-grained cues for temporal reasoning within each clip, while our BiLSTM module captures long-range temporal relationships between multiple clips to detect the given event within the video effectively. The lack of a time-aware encoder reduces the performance of LongVLM [49], while TimeChat [35] suffers from a lack of fine-grained information.
@@ 5.2 Supervised Results
We now evaluate all methods in the supervised setting.
5.2.1 Dense Video Captioning, Video Highlight Detection, and Temporal Video Grounding Results
@@
We present the results of all methods in the supervised setting in Tab. 2. It can be seen from Tab. 2 that our TemporalVLM model achieves the best results across all tasks and metrics. As compared to the zero-shot results in Tab. 1, we see a further increase in performance, e.g., +24.3 on R@1 (IoU=0.5) in the temporal video grounding task and +9.5 on CIDEr in the dense video captioning task. However, we acknowledge that task-specific models such as MMN [48] and QD-DETR [29] may outperform generalist models such as TimeChat [35] and our TemporalVLM model due to their task-specific design paradigms such as using a mutually matching network to directly model the relationship between the query text and video content and introducing specialized tokens and loss functions during fine-tuning.
5.2.2 Temporal Action Segmentation Results
@@
We further test all methods on our IndustryASM dataset for temporal action segmentation. Since the activities in IndustryASM significantly differ from those in TimeIT [35] and Valley [27], we fine-tune all models on IndustryASM before evaluation. Tab. 3 presents the results. It is evident from Tab. 3 that TemporalVLM performs the best across all metrics. For example, it outperforms TimeChat [35] by +23.5 and +6.9 on Edit and Acc respectively, and LongVLM [49] by +8.9 and +5.3 on Edit and Acc respectively. The results confirm the benefits of our time-aware and multi-level features for temporal action segmentation.
@@ 5.3 Ablation Results
# Clips SODA_c CIDEr F1
2 2.6 9.2 18.5
4 3.1 11.3 18.9
6 3.4 13.2 20.0
@@
Below we study the effects of the number of short-term clips and BiLSTM. For this purpose, we fine-tune multiple configurations of TemporalVLM on YouCook2 [58].
5.3.1 Impacts of the Number of Short-Term Clips
5.3.2 Impacts of the BiLSTM Module
To study the effects of BiLSTM, we replace it with various alternatives and report the results in Tab. 5. From Tab. 5, BiLSTM outperforms average pooling, which is not trained and employed in prior works [28, 49], and traditional LSTM, which only performs the forward pass and hence relies only on past cues. In contrast, BiLSTM is learnable and conducts both forward and backward passes to utilize both past and future cues. Moreover, we replace BiLSTM with linear layer to examine the impacts of video Q-Formers [56, 35] alone, leading to worse results. This highlights the importance of BiLSTM even when video Q-Formers [56, 35] are being used. Finally, BiLSTM outperforms transformer [43], which we will discuss below.

### v4

@@ Abstract
This paper introduces TemporalVLM, a video large language model (video LLM) capable of effective temporal reasoning and fine-grained understanding in long videos. At the core, our approach includes a visual encoder for mapping a long-term input video into features which are time-aware and contain both local and global cues. In particular, it first divides the input video into short-term clips, which are jointly encoded with their timestamps and fused across overlapping temporal windows into time-sensitive local features. Next, the local features are passed through a bidirectional long short-term memory (BiLSTM) module for global feature aggregation. The extracted time-aware and multi-level features are important for accurate temporal reasoning and fine-grained understanding in long videos. Moreover, to facilitate the evaluation of TemporalVLM, we present a large-scale long video dataset of industry assembly processes, namely IndustryASM, which consists of videos recorded on factory floors with actions and timestamps annotated by industrial engineers for time and motion studies and temporal action segmentation evaluation. Finally, extensive experiments on datasets of long videos, including TimeIT and IndustryASM, show that TemporalVLM achieves superior performance than previous methods across temporal reasoning and fine-grained understanding tasks, namely dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation. To the best of our knowledge, our work is the first to incorporate LSTMs into video LLMs.
@@ 1 Introduction
We propose TemporalVLM, a video LLM capable of effective temporal reasoning and fine-grained understanding in long videos. Fig. 1 shows the differences in model architectures of TemporalVLM and previous works. Our model includes two key architectural contributions for temporal reasoning in long videos. Firstly, we divide a long-term input video into multiple short-term clips and introduce a time-sensitive clip encoder for extracting fused time-aware local features from each clip. Secondly, we adopt a bidirectional long short-term memory (BiLSTM) module which takes all the local features as inputs and computes global features from multiple clips. As a consequence, our features not only are time-sensitive but also contain both local fine-grained and global semantic information, which are crucial for temporal reasoning in long videos. Moreover, to further evaluate our model, we present IndustryASM, a large-scale long video dataset of industry assembly processes for temporal action segmentation benchmarking and time and motion studies. Our IndustryASM dataset comprises of $4803$ videos with an average video duration of $119$ seconds. It covers in total $52$ diverse industry assembly tasks and includes timestamp and action labels. We convert the labels into chat samples with manually written instructions. To the best of our knowledge, it is the first large-scale video dataset recorded on factory floors and labeled by industrial engineers for temporal action segmentation evaluation and time and motion analyses. Lastly, extensive experiments on datasets of long videos, i.e., TimeIT [36] and IndustryASM, demonstrate that our TemporalVLM approach achieves superior results over previous methods across temporal reasoning and fine-grained understanding tasks, namely dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation.
@@
  • We develop TemporalVLM, a video LLM for temporal reasoning and fine-grained understanding in long videos. Given a long-term input video, we first sample short-term clips and encode each clip along with its timestamps into fused time-aware local features. The local features are then fed to a BiLSTM module to extract global features. Therefore, our features not only are time-sensitive but also capture effective local and global information.
@@ 3 TemporalVLM: A Video LLM for Temporal Reasoning in Long Videos
Figure 2: Given a long-term input video with timestamps, we divide it into short-term clips with timestamps. The time-aware clip encoder is employed on each clip along with its timestamps for extracting fused time-aware local features, which capture local fine-grained cues within each clip. Next, the bidirectional long short-term memory module is adopted for computing global features, which capture global temporal relationships across multiple clips. Lastly, the large language model takes video and text tokens as inputs and generates responses.
@@ 3.1 Time-Aware Clip Encoder
We propose a time-aware clip encoder (i.e., blue block in Fig. 2) for extracting local fine-grained cues within a clip. Given a long-term input video, we first decompose it into $C$ short-term clips and uniformly sample a fixed number of frames $N^{c}_{f}$ from each clip. Here, we set $C=6$ clips and $N^{c}_{f}=96$ frames per clip. The sampled frames for each clip along with their corresponding timestamps are then passed to our time-aware clip encoder for simultaneously encoding the frame contents and their timestamps, yielding fused time-aware local features. The time-aware frame encoder uses a pre-trained image encoder [40] to obtain frame features, which are jointly encoded with timestamps via an image Q-Former [5], yielding time-aware frame features $\mathbf{f}_{t}$. These frame features $\mathbf{f}_{t}$ are then fed to a video Q-Former  [36] across overlapping windows $\mathbf{W}_{i}$ of size $q$ = 32 and overlap $o$ = 16 for temporal modeling. The window-wise outputs $\mathbf{V}_{i}$ are then concatenated into $\mathbf{S}$ :
[…]
$\mathbf{S}$ $=[\mathbf{V}_{1},\mathbf{V}_{2},\dots,\mathbf{V}_{n}],$ (2)
with $n$ denoting the number of overlapping windows within a short-term clip. We propose a fusion module to align diverse temporal cues from redundant boundary tokens in $\mathbf{S}$. Specifically, we apply multiheaded self-attention on $\mathbf{S}$, yielding $\mathbf{C}$ which fuses local contexts across multiple windows into a single context-aware embedding as:
$\mathbf{C}^{(h)}$ $=\text{SoftMax}\left(\frac{\mathbf{Q}^{(h)}{\mathbf{K}^{(h)}}}{\sqrt{d}}\right)\mathbf{V}^{(h)},$ (3)
$\mathbf{C}$ $=[\mathbf{C}^{(1)},\mathbf{C}^{(2)},\dots,\mathbf{C}^{(H)}]\mathbf{W}^{O}$ (4)
We first project $\mathbf{S}$ into $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$ for each attention head $h$. At each head $h$, $\mathbf{C}^{(h)}$ represents the dot product attention between queries $\mathbf{Q}^{(h)}$ and keys $\mathbf{K}^{(h)}$ scaled over dimension $d$ followed by softmax weighted aggregation over $\mathbf{V}^{(h)}$. $\mathbf{C}$ contains the aggregated outputs from all heads with final projection $\mathbf{W}^{O}$ applied. Using overlapping windows leads to spatially redundant time-aware tokens in $\mathbf{S}$, each with different local window context. Through Eq. 4, each redundant time-aware token in $\mathbf{S}$ is enabled to compare and weigh its contextually dissimilar variant, aligning them into a fused temporal representation. This allows $\mathbf{C}$ to yield rich time-aware clip level information by leveraging diverse temporal views from across overlapping windows. Recent works, e.g., [50], leverage clip sampler and clip encoder for long video understanding but their encoders are not timeaware and hence are not effective for temporal reasoning.
@@ 3.2 Bidirectional Long Short-Term Memory
We introduce a BiLSTM module (i.e., green block in Fig. 2) for computing global features across multiple clips. We first concatenate time-aware local features extracted from clips in the temporal order which the clips appear in the video. We then pass the sequence of fused local features to a BiLSTM module for aggregating global features. Our BiLSTM module follows a standard architecture, including two LSTM networks: one processes the sequence in the original order (forward LSTM) and another processes the sequence in the reverse order (backward LSTM) as:
@@
where $\mathbf{h}^{f}_{t}$ and $\mathbf{h}^{b}_{t}$ denote the hidden states at time step $t$ of the forward LSTM and backward LSTM respectively and $\mathbf{c}_{t}$ is the input at time step $t$. The final output $\mathbf{h}_{t}$ at time step $t$ is obtained by concatenating the outputs of the forward LSTM and backward LSTM as $\mathbf{h}_{t}=[\mathbf{h}^{f}_{t},\mathbf{h}^{b}_{t}]$. Please refer to the textbook [19] for a detailed description. Our BiLSTM module utilizes both past and future information and is capable of capturing long-range temporal relationships in both forward and backward directions. As empirically observed in Sec. 5.3, BiLSTM outperforms various alternatives, including average pooling, linear layer, traditional LSTM, and transformer [44]. Our final features output by BiLSTM not only are time-sensitive but also contain both local fine-grained and global semantic information, which are important for temporal reasoning in long videos. Unlike LongVLM [50], which applies pooling operation to obtain global features, our BiLSTM module has learnable parameters and achieves better results as shown in Sec. 5.
@@ 4 IndustryASM: A Large-Scale Long Video Dataset of Industry Assembly Processes
We introduce IndustryASM, a large-scale long video dataset of industry assembly processes for temporal action segmentation and time and motion studies.
@@ 5 Experiments
Implementation Details. We follow TimeChat [36] to use LLaMA-2 7B [43] as the LLM, and ViT-G/14 from EVA-CLIP [40] as the image encoder. We take InstructBLIP’s checkpoints to initialize the image Q-Former, and Video-LLaMA’s [57] checkpoints to initialize the video Q-Former. We employ pytorch’s [15] implementation of multiheaded attention in the fusion module of the clip encoder with 8 attention heads. The BiLSTM module includes a forward LSTM with one hidden layer and a backward LSTM with one hidden layer. The parameters of the BiLSTM module and projection layers are initialized randomly. The LLM is fine-tuned using LoRA [12] with rank 32. We use 6 clips and sample 96 frames from each clip. We fine-tune our model on a subset of TimeIT [36] and Valley [28] on a single 8-A100 (80 GB) machine. The parameters of the LLM and image encoder remain frozen throughout fine-tuning. Please see our supplementary material for more details.
@@ 5.1 Zero-Shot Results
Dense Video Captioning Results. Tab. 1 shows that our TemporalVLM model achieves the best results across all three metrics, i.e., SODA_c, CIDEr, and F1 Score. For example, we outperform TimeChat [36] by +0.3 on CIDEr and +0.5 on F1 Score. This is likely because TimeChat [36] treats the entire video as a single clip and aggregates video tokens via query aggregation, overlooking fine-grained information. In contrast, our extracted features, which capture both local fine-grained cues from each clip and global temporal dependencies across multiple clips, are effective for caption generation and event localization. Moreover, our approach obtains superior results than LongVLM [50] across SODA_c (+0.4), CIDEr (+1.2), and F1 Score (+0.8), which confirms the importance of our time-aware clip encoder and BiLSTM for dense video captioning.
Video Highlight Detection Results. The video highlight detection results in Tab. 1 show that our TemporalVLM model obtains the best performance on both mAP and HIT@1. In particular, it surpasses TimeChat [36] by +1.9 in mAP and +7.4 in HIT@1, while outperforming LongVLM [50] by +5.3 at mAP and by +16.3 at HIT@1. These results suggest the effectiveness of our model for video highlight detection. By aggregating global features from time-aware local features, it identifies the most relevant timestamps and assigns them saliency scores based on their ranks of importance within the video against a given query. Though LongVLM [50] extracts multi-level features, the lack of a time-aware encoder hinders their results.
Temporal Video Grounding Results. It is evident from Tab. 1 that TemporalVLM achieves the best performance for temporal video grounding across R@1 (IoU=0.5) and R@1 (IoU=0.7). Specifically, it outperforms TimeChat [36] by +2.2 and +0.9 on R@1 (IoU=0.5) and R@1 (IoU=0.7) respectively, and LongVLM [50] by +16.2 and +7.1 on R@1 (IoU=0.5) and R@1 (IoU=0.7) respectively. The results validate the benefits of our time-aware and multi-level features for temporal video grounding. Our clip encoder extracts fine-grained cues for temporal reasoning within each clip via fusion, which enables our BiLSTM module to capture long-range temporal relationships between multiple clips to detect the given event effectively. The lack of a time-aware encoder affects LongVLM [50], while TimeChat [36] suffers from a lack of fine-grained information.
@@ 5.2 Supervised Results
Dense Video Captioning, Video Highlight Detection, and Temporal Video Grounding Results. We present the results of all methods in the supervised setting in Tab. 2. Our TemporalVLM model achieves the best results across all tasks and metrics. As compared to the zero-shot results in Tab. 1, we see a further increase in performance, e.g., +24.3 on R@1 (IoU=0.5) in the temporal video grounding task and +9.5 on CIDEr in the dense video captioning task. However, we acknowledge that task-specific models such as MMN [49] and QD-DETR [30] may outperform generalist models such as TimeChat [36] and our TemporalVLM model due to their task-specific design paradigms, e.g., using a mutually matching network to directly model the relationship between the query text and video content or specialized tokens and loss functions during fine-tuning.
@@
Temporal Action Segmentation Results. We further test all methods on our IndustryASM dataset for temporal action segmentation. Since the activities in IndustryASM significantly differ from those in TimeIT [36] and Valley [28], we fine-tune all models on IndustryASM before evaluation. Tab. 3 presents the results. TemporalVLM performs the best across all metrics. For example, it outperforms TimeChat [36] by +23.5 and +6.9 on Edit and Acc respectively, and LongVLM [50] by +8.9 and +5.3 on Edit and Acc respectively. The results confirm the benefits of our fused time-aware and multi-level features.
@@ 5.3 Ablation Results
Below we study the effects of our overlapping sliding video Q-Former and BiLSTM. For this purpose, we fine-tune multiple configurations of TemporalVLM on YouCook2 [59].
Impacts of Overlapping Sliding Video Q-Former. We conduct experiments in Tab. 4 to demonstrate the importance of fusing encoded frames across overlapping temporal windows in our clip encoder. Our model with fusion across overlapping windows achieves the best results, whereas using non-overlapping windows or removing the fusion module leads to performance drops. The results validate the design of our overlapping sliding video Q-Former module.
Model SODA CIDEr F1
No Overlap + No Fusion 2.3 11.0 17.3
No Overlap + Fusion 2.0 9.5 16.2
Overlap + No Fusion 2.8 11.7 18.9
Overlap + Fusion 3.4 13.2 20.0
Table 4: Effects of overlapping sliding video Q-Former. Best results are in bold, second best are underlined.
Impacts of BiLSTM. To study the effects of BiLSTM, we replace it with various alternatives and report the results in Tab. 5. From Tab. 5, BiLSTM outperforms average pooling, which is not trained and employed in prior works [29, 50], and traditional LSTM, which only performs the forward pass and hence relies only on past cues. In contrast, BiLSTM is learnable and conducts both forward and backward passes to utilize both past and future cues. Moreover, we replace BiLSTM with linear layer to examine the impacts of video Q-Formers [57, 36] alone, leading to worse results. This highlights the importance of BiLSTM even when video Q-Formers [57, 36] are being used. Finally, BiLSTM outperforms transformer [44], which we will discuss below.